I need to find the center and radius of a circle with the given formula:
x²+y²=36

Answer by stanbon(75887)

I need to find the center and radius of a circle with the given formula:
x²+y²=36
----
Rewrite in center-radius form:
(x-0)^2 + (y-0)^2 = 6^2
------
Center: (0,0)
Radius: 6
=============================
Cheers,
Stan H.
=============